The Corporation has two production departments, Casting and Finishing. The company uses a job-order costing system and computes a predetermined overhead rate in each production department. The Casting Department's predetermined overhead rate is based on machine-hours and the Finishing Department's predetermined overhead rate is based on direct labor-hours. At the beginning of the current year, the company had made the following estimates:

Casting

Finishing

Machine-hours

17,000

7,000

Direct labor-hours

2,000

15,000

Total fixed manufacturing overhead cost

$

105,400

$

52,800

Variable manufacturing overhead per machine-hour

$

1.70

Variable manufacturing overhead per direct labor-hour

$

3.90

Below is actual information related to Job 215 for the period:

                                                                                                Casting            Finishing

Actual machine-hours                                                             325                  40

Actual direct labor-hours                                                        50                    120

Direct materials                                                                       $1,200             $530

Direct labor                                                                             $500                $960

Please compute the total product cost of Job 215.
Answer #1

Casting overhead rate = (105400/17000+1.7) = 7.9 per MH

Finishing overhead rate = (52800/15000)+3.9 = 7.42 per DLH

Calculate total product cost

Direct material (1200+530) 1730
Direct labor (500+960) 1460
Casting overhead (325*7.9) 2567.50
Finishing overhead (120*7.42) 890.40
Total Product cost 6647.90 or 6648